What is the linear equation that has the following solutions?
x= -9 y= 1856
x= -6 y= 1226

Answer by macston(5194)

.
Given two points: (-9,1856) and (-6,1226)
.
Find the slope (m)=Change in y/change in x
.
m=(1856-1226)/-9-(-6)=630/-3=-210
The slope is -210. Use this for m in slope intercept form:
.
y=mx+b
y=-210x+b Put in values from one point and solve for b.
1856=-210(-9)+b
1856=1890+b
-34=b
The y intercept (b)=-34 Use this as b in slope-intercept form:
.
y=-210x-34 ANSWER: This is your equation. Use second point to check:
.
CHECK:
Using (-6,1226):
1226=-210(-6)-34
1226=1260-34
1226=1226